# The Radio Counterpart to the Fast X-ray Transient EP240414a

### v1

@@ 4.3.1 Low Bulk Velocity
for a completely filled emitting region in the Newtonian limit. The distance is given in units of $10^{28}\,\rm{cm}$, the flux in mJy, the frequency in units of $10\,\rm{GHz}$, and $z$ is the redshift of the source. Using the peak of our $3\,\rm{GHz}$ light curve this implies $E_{\rm{eq}}=1.5\times 10^{49}\,\rm{erg}$ and $R_{\rm{eq}}=1.8\times 10^{17}\,\rm{cm}$. At 20 days post explosion (in the source rest frame) this size implies an expansion velocity $\beta_{\rm{eq,N}}\approx 3.3c$. Note that while Matsumoto & Piran (2023) assumes a different geometry to Fender & Bright (2019) (a conical outflow and an expanding sphere, respectively) the derived radii and energies agree to within a factor of order unity. This result implies that the source has a significant, likely at least mildly relativistic, expansion velocity.
@@ 4.3.2 Relativistic Considerations
Figure 4: Solution for the minimum bulk Lorentz factor for Equation 4 and the one given in Matsumoto & Piran (2023), which becomes significantly incorrect for $\beta_{\rm eq,N}\lesssim 15$ at which point the ratio becomes larger than 1.05. At $\beta_{\rm eq,N}\lesssim 9$ the ratio surpasses 1.1.
@@
Our results imply that the emitting region responsible for the radio emission from EP240414a was expanding at least moderately relativistically at 30 days post explosion. The Newtonian equipartition energy ($E_{\rm{eq},\rm{N}}\approx 1.5\times 10^{49}\,\rm{erg}$) is reduced by a factor of $\sim 10$ for the on-axis relativistic case (e.g. $e_{\rm{on}}=0.1$). Such constraints are similar to those derived from equipartition in other long GRBs (e.g. Bright et al., 2023), however the potentially large Lorentz factors and uncertain geometries of GRB jets make comparison challenging, for example if the opening angle of the jet $\theta_{j}>1/\Gamma$ then the energy is underestimated by a factor or $4\Gamma^{2}(1-\cos\theta_{j})$.
@@ 5 Conclusions
We present radio observations of EP240414a, the second FXT with a radio counterpart. The rise time and peak luminosity of EP240414a distinguish it from regular supernovae and tidal disruption events but are commonly seen from gamma-ray bursts and perhaps from the most luminous fast blue optical transients. Due to the association of EP240414a with a Ic broad line supernova (van Dalen et al. 2024) a short gamma-ray burst is disfavoured despite the large spatial offset from the host galaxy. Based on minimum energy arguments and the presence of a turn-over in our $3\,\rm{GHz}$ light curve of EP2402414a we place approximate limits on the bulk Lorentz factor of the outflow of $\Gamma\gtrsim 1.6$ with an energy $\sim 10^{48}\,\rm{erg}$ in the on-axis case. Additionally, we extend the results presented in Matsumoto & Piran (2023) to show that a general minimum energy solution for on-axis ($\theta\rightarrow 0$) sources exists for any measured apparent source velocity, without the assumption of a large bulk Lorentz factor (Equation 4). Our observations suggest that at least a moderately relativistic outflow was present in EP240414a and that the progenitor and engine was likely a collapsar, such as those that produce the known population of gamma ray bursts. The absence of a reported gamma-ray counterpart to EP240414a suggests that it could be a GRB with a relatively low isotropic equivalent gamma-ray energy (below $\sim 10^{51}\,\rm{erg}$), or that the radio emission is driven by a mildly off-axis jet (as as been suggested by e.g. Ibrahimzade et al., 2024; Wichern et al., 2024) or a cocoon associated with a choked jet (e.g. Bromberg et al., 2012). Such scenarios have been suggested for relativistic supernovae such as SN2009bb (e.g. Soderberg et al., 2010), however EP240414a is significantly more luminous than such events. Based on the observations presented in this work, Gillanders et al. (2024), and Ricci et al. (2024), we suggest that a significant fraction of EP FXTs are associated with the collapse of massive stars and their afterglow.

### v2

@@ 4.3.1 Low Bulk Velocity
for a completely filled emitting region in the Newtonian limit. The distance to the source ($d_{L,28}$) is given in units of $10^{28}\,\rm{cm}$, the flux density ($F_{\nu,\rm{mJy}}$) in mJy, the frequency ($\nu$) in units of $10\,\rm{GHz}$, and $z$ is the redshift of the source. Using the peak of our $3\,\rm{GHz}$ light curve this implies $E_{\rm{eq}}=1.5\times 10^{49}\,\rm{erg}$ and $R_{\rm{eq}}=1.8\times 10^{17}\,\rm{cm}$. At 20 days post explosion (in the source rest frame) this size implies an expansion velocity $\beta_{\rm{eq,N}}\approx 3.3c$. Note that while Matsumoto & Piran (2023) assumes a different geometry to Fender & Bright (2019) (a conical outflow and an expanding sphere, respectively) the derived radii and energies agree to within a factor of order unity. This result implies that the source has a significant, likely at least mildly relativistic, expansion velocity.
@@ 4.3.2 Relativistic Considerations
Figure 4: Solution for the minimum bulk Lorentz factor for Equation 4 and the one given in Matsumoto & Piran (2023), which becomes significantly incorrect for $\beta_{\rm eq,N}\lesssim 15$ at which point the ratio becomes larger than 1.05. At $\beta_{\rm eq,N}\lesssim 9$ the ratio surpasses $1.$
@@
Our results imply that the emitting region responsible for the radio emission from EP240414a was expanding at least moderately relativistically at 30 days post explosion. The Newtonian equipartition energy ($E_{\rm{eq},\rm{N}}\approx 1.5\times 10^{49}\,\rm{erg}$) is reduced by a factor of $\sim 10$ for the on-axis relativistic case (e.g. $e_{\rm{on}}=0.1$). The potentially large Lorentz factors and uncertain geometries of GRB jets make comparison challenging, for example if the opening angle of the jet $\theta_{j}>1/\Gamma$ then the energy is underestimated by a factor or $4\Gamma^{2}(1-\cos\theta_{j})$.
@@ 5 Conclusions
We present radio observations of EP240414a, the second FXT with a radio counterpart. The rise time and peak luminosity of EP240414a distinguish it from regular supernovae and tidal disruption events but are commonly seen from gamma-ray bursts and perhaps from the most luminous fast blue optical transients. Due to the association of EP240414a with a Ic broad line supernova (van Dalen et al., 2024) a short gamma-ray burst is disfavoured despite the large spatial offset from the host galaxy. Based on minimum energy arguments and the presence of a turn-over in our $3\,\rm{GHz}$ light curve of EP2402414a we place approximate limits on the bulk Lorentz factor of the outflow of $\Gamma\gtrsim 1.6$ with an energy $\sim 10^{48}\,\rm{erg}$ in the on-axis case. Additionally, we extend the results presented in Matsumoto & Piran (2023) to show that a general minimum energy solution for on-axis ($\theta\rightarrow 0$) sources exists for any measured apparent source velocity, without the assumption of a large bulk Lorentz factor (Equation 4). Our observations suggest that at least a moderately relativistic outflow was present in EP240414a and that the progenitor and engine was likely a collapsar, such as those that produce the known population of gamma ray bursts. The absence of a reported gamma-ray counterpart to EP240414a suggests that it could be a GRB with a relatively low isotropic equivalent gamma-ray energy (below $\sim 10^{51}\,\rm{erg}$), or that the radio emission is driven by a mildly off-axis jet (as as been suggested by e.g. Ibrahimzade et al., 2024; Wichern et al., 2024) or a cocoon associated with a choked jet (e.g. Bromberg et al., 2012). Such scenarios have been suggested for relativistic supernovae such as SN2009bb (e.g. Soderberg et al., 2010), however EP240414a is significantly more luminous than such events. Based on the observations presented in this work, Gillanders et al. (2024), and Ricci et al. (2024), we suggest that a significant fraction of EP FXTs are associated with the collapse of massive stars and their afterglow.